# Grading Massive Open Online Courses Using Large Language Models

### v1

@@ 3.3 Baseline
We analyze average scores produced by the LLMs for each question across all courses, comparing them to the peer grading as the baseline and the instructor-assigned grades as the ground truth. By evaluating the average scores with respect to the ground truth, our goal is to identify the most effective type of prompt and the corresponding LLM. In this context, a prompt, along with the respective LLM, is deemed superior if it demonstrates greater alignment with the instructor-assigned grades compared to those obtained through peer grading.
Table 1: Average grades from grading 10 students’ writing assignments using the three proposed prompting techniques: ZCoT with instructor-provided correct answers, ZCoT combined with instructor-provided correct answers and rubrics, and ZCoT with instructor-provided correct answers and LLM-generated rubrics. Results encompass three MOOCs: Introductory Astronomy, Astrobiology, and the History and Philosophy of Astronomy. Grading is conducted using GPT-3.5 and GPT-4 as the underlying LLMs. Each average grade includes a subscripted standard deviation. As the differences in average grades for both peer and LLM grades relative to instructor grades are not statistically significant, no grades are marked with an asterisk (see Appendix B for details on $p$-values).
@@ 4 Experimental Setup
Data. This study employs data from three MOOCs: Astrobiology (Impey, 2024a), Introductory Astronomy (Impey, 2024b), and the History and Philosophy of Astronomy (Impey, 2024c). The process of obtaining the data followed a strict ethical roadmap, guided by an Institutional Review Board from the university. The participants willingly filled out a survey and formally consented to allow their course data to be used for research. All data related to the students was de-identified and purposefully sampled to represent a diverse range of grades.
Although we did our best to avoid bias in the sampling process, we are aware that these courses might not be universally representative of the global population, due to the inherent limitations in subsampling from the entire population. Nevertheless, they are representative of the audience on Coursera [^2] https://www.coursera.org as a whole. Most students in these MOOCs are between the ages of 25 and 44. A slight majority of students are male across all courses, with the smallest gender gap in the Astrobiology course (45% female vs. 54% male) and the largest in the History and Philosophy course (34% female vs. 65% male). Geographically, most students are from the United States and India.
@@
In our experiments, we analyze a subset of 10 student writing assignments per course, with each course containing multiple assignments. Also, note that the data in this study is sourced from a proprietary repository, proactively accounting for the potential issue of data contamination in LLMs Golchin and Surdeanu (2023b, a).
@@ 5 Results and Discussion
Table 1 and Table 2 present the outcomes of our assessment of writing assignments using GPT-4 and GPT-3.5 across three MOOC subjects. These assessments are grounded on three different prompts devised in our study. Table 1 shows the average grades derived directly from grading the assignments of 10 students per question for each course, whereas Table 2 displays the results after conducting bootstrap resampling with 10,000 iterations on the data from Table 1. In the following, we provide both our qualitative and quantitative analyses.
@@ 5.1 Qualitative Analysis: Average Grades Evaluation
(2) Both Table 1 and Table 2 reveal that GPT-4 overall outperforms GPT-3.5 by generating grades that are more closely aligned with the grades provided by instructors, especially in the context of the History and Philosophy of Astronomy. This remains constant across all three prompts assessed.
(3) Focusing on the prompts, ones that integrate both instructor-provided rubrics and LLM-generated rubrics give rise to grades that are better aligned with grades given by instructors. This is true for both GPT-4 and GPT-3.5.
(4) In almost all scenarios, GPT-4, when prompted with instructor-provided correct answers and rubrics, generates grades that are either superior to or on par with peer grades. This even includes the History and Philosophy of Astronomy, which needs speculative or imaginative thinking abilities.
(5) When GPT-4 serves as the underlying base model, LLM-generated rubrics demonstrate comparable performance to instructor-provided rubrics, as illustrated in Table 1 and Table 2. This finding is interesting, indicating the potential use of LLMs for rubric generation in educational contexts, enabling the automation of grading processes with only the input of correct answers.
(6) From the results listed in Table 1 and Table 2, it is clear that the most difficult course to grade is the History and Philosophy of Astronomy. Both LLMs and peer graders struggle to match instructor-assigned grades closely. Nevertheless, even in these cases, GPT-4 outperforms peer grading, generating grades more aligned with instructor-assigned ones.
@@ 5.2 Quantitative Analysis: Question-by-Question Evaluation
(1) Generally, LLM-assigned scores are higher and more generous than instructor-assigned grades. Although these scores are higher, they align more closely with instructor-assigned grades than peer grades do, particularly in the History and Philosophy of Astronomy. Despite this alignment, some instability was observed. For example, in the Astrobiology course, both models vary dramatically regardless of the prompt strategy employed, with GPT-3.5 scoring low and GPT-4 scoring high, differing from the instructor-assigned scores. We also observed that LLMs encounter challenges based on the length of students’ answers, particularly with extremely short or excessively long responses. In such cases, while GPT-4 handled short answers well, it struggled with long ones.
@@ 6 Conclusion
In sum, our results show the potential of using LLMs in MOOCs to automate grading systems and replace peer grading, thus enhancing the learning experience for millions of online students globally.
@@ Appendix A Details on Assignment Questions
Appendix B Details on $p$-values
As detailed in Section 4: Bootstrap Resampling, we use 10,000 iterations with replacement in all settings with bootstrap resampling. Table 3 presents the $p$-values for the differences in average scores referenced in Table 1. Note that the significance of average differences is assessed using a threshold of 0.05. Based on the computed $p$-values, none of the differences in average scores is statistically significant.

### v2

@@ 3.3 Baseline
We analyze average scores and MAEs produced by the LLMs for each question across all courses, comparing them to the peer grading as the baseline and the instructor-assigned grades as the ground truth. By evaluating the average scores and MAEs with respect to the ground truth, our goal is to identify the most effective type of prompt and the corresponding LLM. In this context, a prompt, along with the respective LLM, is deemed superior if it demonstrates greater alignment with the instructor-assigned grades compared to those obtained through peer grading.
Table 1: Average grades from grading 10 students’ writing assignments using the three proposed prompting techniques: ZCoT with instructor-provided correct answers, ZCoT combined with instructor-provided correct answers and rubrics, and ZCoT with instructor-provided correct answers and LLM-generated rubrics. Results encompass three MOOCs: Introductory Astronomy, Astrobiology, and the History and Philosophy of Astronomy. Grading is conducted using GPT-3.5 and GPT-4 as the underlying LLMs. Each average grade includes a subscripted standard deviation. As the differences in average grades for both peer and LLM grades relative to instructor grades are not statistically significant, no grades are marked with an asterisk (see Appendix C for details on $p$-values).
@@ 4 Experimental Setup
Data. This study employs data from three MOOCs: Astrobiology (Impey, 2024a), Introductory Astronomy (Impey, 2024b), and the History and Philosophy of Astronomy (Impey, 2024c). The process of obtaining the data followed a strict ethical roadmap, guided by an Institutional Review Board from the university. Participants voluntarily completed a survey and provided formal consent for their data to be used in the research. Also, all identifying information was anonymized. Since the original dataset had most assignments clustered at either very high or low grades, we sampled data to cover a wider range of grades for a more comprehensive evaluation of LLM grading performance.
Most students in these MOOCs are between the ages of 25 and 44. A slight majority of students are male across all courses, with the smallest gender gap in the Astrobiology course (45% female vs. 54% male) and the largest in the History and Philosophy course (34% female vs. 65% male). Geographically, most students are from the United States and India.
@@
Following previous studies (Pankiewicz and Baker, 2023; Kiesler et al., 2023; Gabbay and Cohen, 2024, inter alia), we use a subset of data for our experiments. Specifically, we analyze 12 assignments from a group of 10 students in each of the three courses, resulting in analyzing 120 questions per prompt. Note that the data in this study is sourced from a proprietary repository, proactively accounting for the potential issue of data contamination in LLMs Golchin and Surdeanu (2023b, a).
@@ 5 Results and Discussion
Tables 1 and 2 present the outcomes of our assessment of writing assignments using GPT-4 and GPT-3.5 across three MOOC subjects. These assessments are grounded on three different prompts devised in our study. Table 1 shows the average grades derived directly from grading the assignments of 10 students per question for each course, whereas Table 2 displays the results after conducting bootstrap resampling with 10,000 iterations on the data from Table 1. In the following, we provide both our qualitative and quantitative analyses.
@@ 5.1 Qualitative Analysis: Average Grades Evaluation
(2) Tables 1 and 2 reveal that GPT-4 overall outperforms GPT-3.5 by generating grades that are more closely aligned with the instructor-assigned grades, especially in the context of the History and Philosophy of Astronomy.
(3) Focusing on the prompts in Tables 1 and 2, ones that integrate both instructor-provided rubrics and LLM-generated rubrics give rise to grades that are better aligned with those from the instructors. This is true for both GPT-4 and GPT-3.5.
(4) Looking at Tables 1 and 2, in almost all scenarios, GPT-4, when prompted with instructor-provided correct answers and rubrics, generates grades that are either superior to or on par with peer grades. This even includes the History and Philosophy of Astronomy, which needs speculative or imaginative thinking abilities.
(5) When GPT-4 serves as the underlying base model, LLM-generated rubrics demonstrate comparable performance to instructor-provided rubrics, as illustrated in Tables 1 and 2. This finding is interesting, indicating the potential use of LLMs for rubric generation in educational contexts, enabling the automation of grading processes with only the input of correct answers.
(6) Tables 3 and 4 (in Appendix B) show that LLMs more often produce the lowest MAEs across all questions compared to peer grading, indicating that LLM-assigned grades are more accurate, with fewer errors or deviations from the instructor-assigned grades. Additionally, GPT-4 consistently assigns grades with lower MAEs than GPT-3.5.
(7) From the results listed in Tables 1 and 2, it is clear that the most difficult course to grade is the History and Philosophy of Astronomy. Both LLMs and peer graders struggle to match instructor-assigned grades closely. Nevertheless, even in these cases, GPT-4 outperforms peer grading, generating grades more aligned with instructor-assigned ones.
@@ 5.2 Quantitative Analysis: Question-by-Question Evaluation
(1) Generally, LLM-assigned grades are higher and more generous than instructor-assigned grades. Although these scores are higher, they align more closely with instructor-assigned grades than peer grades do, particularly in the History and Philosophy of Astronomy. Despite this alignment, some instability was observed. For example, in the Astrobiology course, both models vary dramatically regardless of the prompt strategy employed, with GPT-3.5 scoring low and GPT-4 scoring high, differing from the instructor-assigned scores. We also observed that LLMs encounter challenges based on the length of students’ answers, particularly with extremely short or excessively long responses. In such cases, while GPT-4 handled short answers well, it struggled with long ones.
@@
(4) In our experiments with GPT-3.5 and GPT-4, we did not observe any instances of hallucinations when generating grades and feedback for assignments. This can be attributed to the inclusion of detailed rubrics and correct answers in the input prompt, which guide the output into a more structured and factual format. These findings suggest the potential of LLMs to serve as reliable automated grading systems when provided with sufficient guidance in the input prompt.
@@ 6 Conclusion
Overall, our results show the potential of using LLMs in MOOCs to automate grading systems and replace peer grading, thus enhancing the learning experience for millions of online students globally.
@@ 7 Limitations
8 Ethical Considerations
The use of LLMs for automated grading introduces several ethical concerns that warrant careful examination. One major concern is fairness in grading. In fact, biases present in the training data of LLMs could lead to systematic disparities in scores across different demographic groups. For example, if an LLM has been trained on data that reflects societal biases, it might inadvertently favor certain linguistic styles, cultural norms, argumentation styles, or perspectives, thereby disadvantaging students who do not conform to these patterns.
Students’ perceptions of machine-generated feedback also present ethical concerns. Feedback from LLMs might be perceived as less personal or less tailored to individual learning paths compared to human-generated feedback. It is essential to communicate clearly with students about how automated grading systems using LLMs operate, the rationale behind their feedback, and the ways in which human oversight is incorporated into the grading process.
In conclusion, while automated grading with LLMs offers significant benefits in efficiency and scalability, it is imperative to address the aforesaid ethical considerations proactively. Ensuring fairness, minimizing biases, and fostering positive perceptions among students are crucial steps toward leveraging the advantages of LLMs while upholding the integrity of the educational experience.
@@ Appendix A Details on Assignment Questions
Appendix B Mean Absolute Error Results
Tables 3 and 4 present the MAE values for LLM-assigned grades and peer grades relative to instructor-assigned grades. The results show that LLMs generally achieve lower MAE values than peer grades, with GPT-4 most frequently obtaining the lowest MAEs.
Table 3: MAE values for 10 students’ writing assignments, comparing LLM-assigned and peer grades to instructor-assigned grades. Bold values indicate the lowest MAE for each question.
GPT-3.5 GPT-4 Courses Questions Peer Grades      ZCoT W/ Answers ZCoT W/ Answers & Rubrics ZCoT W/ Answers & LLM Rubrics      ZCoT W/ Answers ZCoT W/ Answers & Rubrics ZCoT W/ Answers & LLM Rubrics Introductory Astronomy Q1 1.85      1.20 1.40 1.60      1.05 1.30 1.30 Q2 1.25      0.80 0.90 1.00      0.45 0.50 0.50 Q3 0.70      1.10 0.90 1.30      0.50 0.40 0.40 Q4 1.15      1.00 0.70 0.80      0.50 0.70 0.55 Q5 2.10      1.30 2.30 1.10      1.30 1.40 1.25 Astrobiology Q1 1.30      1.60 1.30 1.68      1.70 0.70 1.30 Q2 1.35      1.90 1.00 2.38      1.40 1.10 1.20 Q3 1.85      2.20 2.00 3.02      1.60 1.80 2.00 History & Philosophy of Astronomy Q1 0.70      0.80 1.50 0.95      0.40 0.50 0.40 Q2 1.30      0.90 1.20 1.04      1.05 1.10 0.75 Q3 1.30      0.80 1.50 1.00      0.95 0.70 0.70 Q4 1.40      0.60 1.30 0.60      1.25 0.90 0.95 Overall Average All 1.35      1.18 1.33 1.37      1.01 0.92 0.94
Table 4: MAE values for 10 students’ writing assignments using bootstrap resampled results with 10,000 iterations, comparing LLM-assigned and peer grades to instructor-assigned grades. Bold values indicate the lowest MAE for each question.
GPT-3.5 GPT-4 Courses Questions Peer Grades      ZCoT W/ Answers ZCoT W/ Answers & Rubrics ZCoT W/ Answers & LLM Rubrics      ZCoT W/ Answers ZCoT W/ Answers & Rubrics ZCoT W/ Answers & LLM Rubrics Introductory Astronomy Q1 1.67      1.30 2.32 2.06      1.72 1.68 1.77 Q2 1.68      1.32 1.01 0.96      0.91 1.02 0.98 Q3 2.43      2.69 2.58 2.85      2.37 2.51 2.35 Q4 2.14      2.46 2.62 2.74      2.34 2.68 2.62 Q5 3.30      2.96 3.55 3.59      3.22 3.53 3.61 Astrobiology Q1 3.35      3.39 3.92 3.62      3.30 3.57 3.47 Q2 3.40      2.93 3.03 4.08      2.93 3.07 3.27 Q3 2.32      3.06 3.27 4.57      2.25 2.51 2.49 History & Philosophy of Astronomy Q1 0.64      0.94 1.50 0.87      0.66 0.86 0.78 Q2 1.44      0.97 1.14 0.94      1.16 0.96 1.00 Q3 1.28      0.82 1.74 1.32      1.02 0.75 0.77 Q4 1.48      0.68 1.42 0.56      1.21 1.00 0.99 Overall Average All 2.09      1.96 2.34 2.35      1.92 2.01 2.01
@@ Appendix B Mean Absolute Error Results
Appendix C Details on $p$-values
As detailed in Section 4 under Bootstrap Resampling, we use 10,000 iterations with replacement in all settings with bootstrap resampling. Table 5 presents the $p$-values for the differences in average scores referenced in Table 1. Note that the significance of average differences is assessed using a threshold of 0.05. Based on the computed $p$-values, none of the differences in average scores is statistically significant.